A 30kg mixture contains pretzels and cheezies. Pretzels are priced at $2.00/kg and cheezies are priced at $5.00/kg. If the mixture is priced at $4.00/kg, how many kgs of each does it contain?
Answer by checkley71(8403)

30*4=X*2+(30-X)*5
120=2X+150-5X
-2X+5X=150-120
3X=30
X=30/3
X=10 KG OF $2.00 PRETZELS
30-10=20 KG OF $ 5.00 CHEEZIES